# Breaking the Trend: How to Avoid Cherry-Picked Signals

### v2

@@ 4 Description of the empirical analysis
To validate Grebenkov’s theory (Eq. 3), introduced in Sec 3 and, therefore, the diffusive process in Eq. 1 which is assumed for the underlying assets, we conducted empirical backtests of the Agnostic Risk Parity (ARP) strategy introduced by Benichou at al (2017). This strategy constructs a rotationally invariant portfolio that applies trend-following while allocating equal unconditional risk to each eigenvector of the covariance matrix. While rotational invariance is a desirable theoretical property, the equal unconditional risk allocation among eigenvectors is a more desired property for asset manager. Moreover, Valeyre (2024) demonstrated that the ARP portfolio is optimal—achieving the highest Sharpe ratio—under the assumption that the covariance of the $\xi_{i,k}$ terms in Eq. 1 follows a particular random matrix structure, which is considered realistic. Implementing the ARP portfolio is relatively straightforward and consists of normalizing the trend indicators by the inverse square root of the correlation matrix, rather than by the full inverse as in the classical Markowitz framework.
@@ 4.1 Renormalized exponential moving average indicator ($\mathrm{EMA}$)
We used only a normalized version of the $EMA$ to account for the dependence of volatility on both time and instrument. This normalization allows for meaningful comparisons of trend indicators across instruments, time periods, and even different smoothing parameters. At time $t$, the $\mathrm{EMA}$ is a vector of $N$ exponential moving averages of normalized returns, one for each of the $N$ instruments. The $EMA$ of instrument $i$ at time $t+1$ is defined by $\phi$ in Eq. 6, where $r_{i,t+1}$ denotes the return of instrument $i$ at time $t+1$. The parameter $\sigma$ represents the standard deviation of returns used to normalize them, and $\eta$ is the smoothing parameter of the $EMA$ as well as of the return normalization. $\sigma^{2}$ can be interpreted as an exponential moving average of squared returns, and the indicator $\phi$ corresponds to an exponential moving average of returns scaled by their volatility.. The incrementation is on daily basis but it could be generalized to minutes returns.
@@
[…]
The indicator $\mathrm{EMA}\left(150\right)$ is defined by the $\phi_{i,t+1}\left(\eta\right)$ when $\eta=\frac{1}{150}$ is applied to every underlying $i$ and time $t$. We use $\sqrt{\eta}$ in Eq.6 so that the std of $\phi_{i,t+1}\left(\eta\right)$ is 1 theoritically if $r$ are well approximated by unautocorrelated returns so that $\phi$ is a normalized $EMA$ indicator.
@@
$\left\{\begin{array}[]{l}\hat{\mathrm{ARP}}_{t+1}=\left(1-\rho\right)\hat{\mathrm{ARP}}_{t}+\rho\Sigma^{-1}_{t}C^{-0.5}\phi_{t}\left(\eta\right)\\[10.0pt] or\\[10.0pt] \hat{\mathrm{ARP}}_{t+1}=\left(1-\rho\right)\hat{\mathrm{ARP}}_{t}+\rho\Sigma^{-1}_{t}C^{-0.5}\mathrm{MACD}_{t}\left(\eta_{1},\eta_{2},\eta_{3},\omega_{1},\omega_{2},\omega_{3}\right)\\ \end{array}\right.$ (12)
@@ 4.4 Dataset and different simulated parameters
Fig.4 displays the sensitivities of these indicators to past daily returns. $\mathrm{MACD}$ enables to put more weights on very older returns as we expected.
Figure 4: Different sensitivities to past daily returns for indicators ($\mathrm{MACD}$, $\mathrm{EMA}$) all based on 20, 80 and 400 days time scales. we have $\mathrm{ARP}\left(0\times 20,80,0.2\right)$, $\mathrm{ARP}\left(0\times 20,80,0.4\times 400\right)$, $\mathrm{ARP}\left(20,80,0.2\times 400\right)$, $\mathrm{ARP}\left(20,80,0.4\times 400\right)$,$\mathrm{ARP}\left(80\right)$, $\mathrm{ARP}\left(400\right)$
@@ 5 Empirical results
Here we first interpret the simulated empirical Sharpe ratio when applying the $\mathrm{ARP}$ portfolio with the diffent indicators. We find interesting results which challenge traditional recipes. Indeed, and quite surprisingly, the empirical Sharpe ratio closely replicates the theoretical Sharpe ratio predicted by Grebenkov and Serror (2014), whose main result shows that a single $EMA$ indicator is already optimal. Secondly, we attempt to replicate the performance of a simple $EMA$ using a basket of more complex trend indicators. Thirdly, we incorporate the non-linear findings of Schmidhuber (2021), which could challenge our assumptions, and finally, we analyze the empirical correlations between different strategies.
@@ 5.1 Grebenkov’s model empirical validation
Interestingly, the empirical simulation fits pretty well the theorical formula of the Sharpe ratio Eq.3 derived in Grebenkov and Serror (2014) with the following parameters $\lambda=1/180$ and $\beta_{0}=0.12$. The Fig.5 is very impressive and the empricial fits should validate the Gebenkov’s model to describe trends. Moreover the fit appears to be more robust than the variogram (Fig.2) of the Dow Jones Index. That is the main result of the paper.
As a first consequence the parameter of $112$ days ($\eta_{\text{opt}}=\frac{1}{112}$) for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio when not accounting cost of trading which are very small at that trading frequency. Lamperiere et al (2014) found a different result, with a faster optimal parameter at 50 days for the EMA, but very similar Sharpe ratios across EMAs ranging from 1 to 5 months, showing an almost flat curve with no clear optimum. The difference with our results may be explained by the choice made in Lamperiere et al (2014) to determine positions based not on the linear magnitude of the EMA signals, but solely on their sign. Additionally, they did not use the ARP portfolio construction, which involves inverting the square root of the correlation matrix—a method that explains a significant part of the Sharpe ratio, as shown in Benichou at al (2017). For these two reasons, their Sharpe ratio for the post-2000 period was measured at 0.85, lower than our 1.2, and their optimal EMA signal corresponded to a faster timescale than ours.
As a second consequence of the very good fit with the theoritical formula, modeling trend through mean-reversion process using only one relaxation time and not an multi-time scales one appears to be a good solution as the fit is more than correct. That is particulary unexpected as market is known to have a multi time scales property: For example the relaxation of volatility is known to have muti-time scales, investors are expected to have different horizons of time and different horizons of analysis. As a consequence we can wonder wether the usual recipe to take into account of a multitude of different indicators is justified to claim having the most robust and the most refined signal.
@@
We can see based on the Tab 4 in appendix C that empirical Sharpe is $1.24$ for $\mathrm{ARP}\left(120\right)$ with one time scale and $1.18$ for $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ with 3 time scales. $\mathrm{MACD}$ could not be justified as additional time scales does not bring significant improvement. Also Sharpe ratio is not so sensitive to the parameters arround the optimal as expected: Sharpe is $1.25$ for $\mathrm{ARP}\left(100\right)$ and $1.21$ for $\mathrm{ARP}\left(150\right)$.
Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) applied to $\mathrm{EMA}$ as signal v.s. $\eta$ the parameter of the $\mathrm{EMA}$ (Eq.6) and the theoretical equation Eq.3. The theoritical model is fitted with parameters $\beta_{0}=0.12$ and $\lambda=1/180$ so that $\eta_{\text{opt}}=\frac{1}{112}$ Empirical measurements are displayed in the tab 4.
@@ 5.2 Replication of a simple $\mathrm{EMA}$ by a mixture of bollinger bands $\mathrm{BB}$
[…]
@@
5.3 Non-linear findings which challenge the $EMA$
@@
A classical approach to introduce mean-reverting behavior from fundamentalist traders is to assume an attraction toward a fair or efficient price though such a price is unobservable. It can be approximated by a moving average, which effectively induces negative autocorrelation, especially over longer time scales. Zakamuli and Giner (2024) explores this idea, though obtaining statistically significant results remains difficult. By contrast, Schmidhuber (2021)’s approach involves making autoregressive coefficients trend-dependent, enabling a dynamic transition between trend-following governed by trend-followers and mean-reversion regimes governed by fundamentalist traders. Another way to model the influence of fundamental investors could be simply to assume their impact manifests in changes to the trend, i.e., on the noise $\xi_{i,k}$ in Eq. 1. It could be enough to make sure that prices are not totally disconnected from fair and efficient ones which would differ in that case from moving average of prices.
The findings of Schmidhuber (2021) should not be ignored. A deeper investigation is however warranted to better understand their implications for the construction of optimal trend-following portfolios, although this lies beyond the scope of the present simple work and should be the subject of a dedicated study.
@@ 5.4 Correlations between indicators
Fig.7 in the Appendix C shows that the indicators with different parameters give very correlated strategies which confims it makes not so appealing the traditional approach of using a basket of many indicators as signal. $\mathrm{ARP}\left(80\right)$ is correlated to $\mathrm{ARP}\left(150\right)$ with a coefficient of $0.96$. That is nevertheless pretty interesting to see so strong correlations which could be explained by a common factor somewhere, which could be the object of an additional research. Morevever the $\mathrm{ARP}\left(120\right)$ based on a simple $\mathrm{EMA}$ is very close with correlation from 1 to 0.99 to the new indicator $\mathrm{MACD}\left(20,120,0\times 400\right)$ or $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ we introduced based on 3 different times scales to increase weights for very old past returns and decrease weights on very recent ones. So it makes this rafinement not justified again.
@@ 6 Conclusion
Grebenkov’s model for describing trends is empirically validated, as its pretty complicated theoretical formula for determining the Sharpe ratio based on the $\mathrm{EMA}$ parameter fits impressively well with empirical data. The best fit is obtained using the theoretical model parameters $\lambda=1/180$ and $\beta_{0}=0.12$. As a consequence the parameter of $112$ days for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio. It is quite surprising that a single $\mathrm{EMA}$ is optimal for capturing trends, as one would expect different time scales for different types of investors. However, there are likely much shorter time scales, on the order of a few days, but they have no significant impact on a medium-frequency strategy. The conclusion is that using a complex mixture of sophisticated indicators is unnecessary when the $\mathrm{EMA}$ alone provides a perfect fit—proving that simplicity can indeed be beautiful.

### v3

@@ 4 Description of the empirical analysis
To validate Grebenkov’s theory (Eq. 3), introduced in Sec 3 and, therefore, the diffusive process in Eq. 1 which is assumed for the underlying assets, we conducted empirical backtests of the Agnostic Risk Parity (ARP) strategy introduced by Benichou at al (2017). This strategy constructs a rotationally invariant portfolio that applies trend-following while allocating equal unconditional risk to each eigenvector of the covariance matrix. While rotational invariance is a desirable theoretical property, the equal unconditional risk allocation among eigenvectors is a more desired property for asset managers. Moreover, Valeyre (2024) demonstrated that the ARP portfolio is optimal—achieving the highest Sharpe ratio—under the assumption that the covariance of the $\xi_{i,k}$ terms in Eq. 1 follows a particular random matrix structure, which is considered realistic. Implementing the ARP portfolio is relatively straightforward and consists of normalizing the trend indicators by the inverse square root of the correlation matrix, rather than by the full inverse as in the classical Markowitz framework.
@@ 4.1 Renormalized exponential moving average indicator ($\mathrm{EMA}$)
We used only a normalized version of the $EMA$ to account for the dependence of volatility on both time and instrument. This normalization allows for meaningful comparisons of trend indicators across instruments, time periods, and even different smoothing parameters. At time $t$, the $\mathrm{EMA}$ is a vector of $N$ exponential moving averages of normalized returns, one for each of the $N$ instruments. The $EMA$ of instrument $i$ at time $t+1$ is defined by $\phi$ in Eq. 6, where $r_{i,t+1}$ denotes the return of instrument $i$ at time $t+1$. The parameter $\sigma$ represents the standard deviation of returns used to normalize them, and $\eta$ is the smoothing parameter of the $EMA$ as well as of the return normalization. $\sigma^{2}$ can be interpreted as an exponential moving average of squared returns, and the indicator $\phi$ corresponds to an exponential moving average of returns scaled by their volatility. The incrementation is on a daily basis but it could be generalized to minutes returns.
@@
$\phi_{i,t+1}\left(\eta\right)=\left(\ 1-\eta\right)\phi_{i,t}\left(\eta\right)+\sqrt{\eta}\frac{r_{i,t+1}}{\sigma_{i,t}\left(\eta\right)}$ (6)
The indicator $\mathrm{EMA}\left(150\right)$ is defined by the $\phi_{i,t+1}\left(\eta\right)$ when $\eta=\frac{1}{150}$ is applied to every underlying $i$ and time $t$. We use $\sqrt{\eta}$ in Eq.6 so that the std of $\phi_{i,t+1}\left(\eta\right)$ is 1 theoretically if $r$ are well approximated by unautocorrelated returns so that $\phi$ is a normalized $EMA$ indicator.
@@
[…]
@@ 4.4 Dataset and different simulated parameters
Fig.4 displays the sensitivities of these indicators to past daily returns. $\mathrm{MACD}$ enables us to put more weights on very older returns as we expected.
Figure 4: Different sensitivities to past daily returns for indicators ($\mathrm{MACD}$, $\mathrm{EMA}$) all based on 20, 80 and 400 day time scales. we have $\mathrm{ARP}\left(0\times 20,80,0.2\right)$, $\mathrm{ARP}\left(0\times 20,80,0.4\times 400\right)$, $\mathrm{ARP}\left(20,80,0.2\times 400\right)$, $\mathrm{ARP}\left(20,80,0.4\times 400\right)$$\mathrm{ARP}\left(80\right)$, $\mathrm{ARP}\left(400\right)$
@@ 5 Empirical results
Here we first interpret the simulated empirical Sharpe ratio when applying the $\mathrm{ARP}$ portfolio with the different indicators. We find interesting results which challenge traditional recipes. Indeed, and quite surprisingly, the empirical Sharpe ratio closely replicates the theoretical Sharpe ratio predicted by Grebenkov and Serror (2014), whose main result shows that a single $EMA$ indicator is already optimal. Secondly, we attempt to replicate the performance of a simple $EMA$ using a basket of more complex trend indicators. Thirdly, we incorporate the non-linear findings of Schmidhuber (2021), which could challenge our assumptions, and finally, we analyze the empirical correlations between different strategies.
@@ 5.1 Grebenkov’s model empirical validation
Interestingly, the empirical simulation fits pretty well the theorical formula of the Sharpe ratio Eq.3 derived in Grebenkov and Serror (2014) with the following parameters $\lambda=1/180$ and $\beta_{0}=0.12$. The Fig.5 is very impressive and the empricial fits should validate Gebenkov’s model to describe trends. The risk of overfitting is very limited, as the formula was originally derived in 2014, with its parameters fitted solely on a long historical series of the Dow Jones Index. Using an extended dataset including commodities, forex, bonds, and other stock indices, we obtain parameter estimates that are consistent with those in Grebenkov and Serror (2014), even if they differ significantly in value. Moreover, the fit appears to be more robust than that obtained using the variogram method (Fig. 2) applied to the Dow Jones Index, as our fitting is performed on a broader set of asset classes over a more recent period. That is the main result of the paper.
As a first consequence the parameter of $112$ days ($\eta_{\text{opt}}=\frac{1}{112}$) for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio when not accounting cost of trading which are very small at that trading frequency (Tab. 4 displays holding periods slower than 80 days). Lamperiere et al (2014) found a different result, with a faster optimal parameter at 50 days for the EMA, but very similar Sharpe ratios across EMAs ranging from 1 to 5 months, showing an almost flat curve with no clear optimum. The difference with our results may be explained by the choice made in Lamperiere et al (2014) to determine positions based not on the linear magnitude of the EMA signals, but solely on their sign. Additionally, they did not use the ARP portfolio construction, which involves inverting the square root of the correlation matrix—a method that explains a significant part of the Sharpe ratio, as shown in Benichou at al (2017). For these two reasons, their Sharpe ratio for the post-2000 period was measured at 0.85, lower than our 1.2, and their optimal EMA signal corresponded to a faster timescale than ours.
As a second consequence of the very good fit with the theoretical formula, modeling trend through mean-reversion process using only one relaxation time and not an multi-time scales one appears to be a good solution as the fit is more than correct. That is particulary unexpected as market is known to have a multi-time scales property: For example the relaxation of volatility is known to have multi-time scales, investors are expected to have different horizons of time and different horizons of analysis. As a consequence we can wonder whether the usual recipe to take into account a multitude of different indicators is justified to claim having the most robust and the most refined signal. When properly renormalizing the signal, the position, and the portfolio, we do not observe trends reversing over longer horizons (e.g., 12 months) that would indicate mean reversion. Instead, the Grebenkov formula still holds even at horizons of up to 1,000 days. This stands in strong contradiction to the findings of Moskowitz et al (2012), who focused on momentum (MOM) indicators over the period 1965–2009.
@@
We can see based on the Tab 4 in appendix C that empirical Sharpe is $1.24$ for $\mathrm{ARP}\left(120\right)$ with one time scale and $1.18$ for $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ with 3 time scales. $\mathrm{MACD}$ could not be justified as additional time scales do not bring significant improvement. Also Sharpe ratio is not so sensitive to the parameters arround the optimal as expected: Sharpe is $1.25$ for $\mathrm{ARP}\left(100\right)$ and $1.21$ for $\mathrm{ARP}\left(150\right)$.
Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) applied to $\mathrm{EMA}$ as signal v.s. $\eta$ the parameter of the $\mathrm{EMA}$ (Eq.6) and the theoretical equation Eq.3. The theoretical model is fitted with parameters $\beta_{0}=0.12$ and $\lambda=1/180$ so that $\eta_{\text{opt}}=\frac{1}{112}$ . Empirical measurements are displayed in the tab 4.
@@ 5.2 Replication of a simple $\mathrm{EMA}$ by a mixture of Bollinger Bands $\mathrm{BB}$
$\mathrm{SMA}_{t}\left(n\right)=\int_{0}^{\infty}\mathrm{BB}_{t}\left(n,\delta\right)\,d\delta$ (16)
@@
5.3 Non-linear findings which challenge the EMAEMA
@@
A classical approach to introduce mean-reverting behavior from fundamentalist traders is to assume an attraction toward a fair or efficient price though such a price is unobservable. It can be approximated by a moving average, which effectively induces negative autocorrelation, especially over longer time scales. Zakamuli and Giner (2024) explore this idea, though obtaining statistically significant results remains difficult. By contrast, Schmidhuber (2021)’s approach involves making autoregressive coefficients trend-dependent, enabling a dynamic transition between trend-following governed by trend-followers and mean-reversion regimes governed by fundamentalist traders. Another way to model the influence of fundamental investors could be simply to assume their impact manifests in changes to the trend, i.e., on the noise $\xi_{i,k}$ in Eq. 1. It could be enough to make sure that prices are not totally disconnected from fair and efficient ones which would differ in that case from moving average of prices.
The findings of Schmidhuber (2021) should not be ignored. A deeper investigation is, however, warranted to better understand their implications for the construction of optimal trend-following portfolios, although this lies beyond the scope of the present simple work and should be the subject of a dedicated study.
@@ 5.4 Correlations between indicators
Fig.7 in Appendix C shows that the indicators with different parameters give very correlated strategies which confims that the traditional approach of using a basket of many indicators as signal is not so appealing. $\mathrm{ARP}\left(80\right)$ is correlated to $\mathrm{ARP}\left(150\right)$ with a coefficient of $0.96$. That is nevertheless pretty interesting to see such strong correlations which could be explained by a common factor somewhere, which could be the object of an additional research. Morevever the $\mathrm{ARP}\left(120\right)$ based on a simple $\mathrm{EMA}$ is very close with correlation from 1 to 0.99 to the new indicator $\mathrm{MACD}\left(20,120,0\times 400\right)$ or $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ we introduced based on 3 different time scales to increase weights for very old past returns and decrease weights on very recent ones. So it makes this refinement not justified again.
@@ 6 Conclusion
Grebenkov’s model for describing trends is empirically validated, as its pretty complicated theoretical formula for determining the Sharpe ratio based on the $\mathrm{EMA}$ parameter fits impressively well with empirical data. The best fit is obtained using the theoretical model parameters $\lambda=1/180$ and $\beta_{0}=0.12$. As a consequence, the parameter of $112$ days for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio. It is quite surprising that a single $\mathrm{EMA}$ is optimal for capturing trends, as one would expect different time scales for different types of investors. However, there are likely much shorter time scales, on the order of a few days, but they have no significant impact on a medium-frequency strategy. The conclusion is that using a complex mixture of sophisticated indicators is unnecessary when the $\mathrm{EMA}$ alone provides a perfect fit—proving that simplicity can indeed be beautiful.
7 Declaration of funding
No funding was received
8 Disclosure of interest
There are no interests to declare,